Mrs. Kim is four years younger than her husband. Find the age of each if the total oh their ages, including that of their one year old is 63.
Answer by mathman409(67)

Lets do this question by using algebra.
First, make the unknown things variables.
Make Mrs. kim "m" and her husband "h".

So, state the knowledge that u have so far.
m=h-4
m+h=62
Now, replace m with h-4

So, h-4+h=62
Simplify
h-4+h=62
2h-4=62
add both sides by 4
2h=66
divide both sides by 2
h=33
now, replace h with 33 to find out m
33-4=m
29=m
now u have the answer:

m=29
h=33

check:
29=33-4
29=29 CORRECT!!!
so, here is the answer once again:
m=29
h=33